OpenStudy (anonymous):

Find the real numbers x and y that make the equation true. 5 + yi = x + 3i

jimthompson5910 (jim_thompson5910):

hint: equate real and imaginary parts together

OpenStudy (anonymous):

For example sub in two numbers, say x=2, y=7 5+yi=x+3i will become 5+7i=2+3i these clearly aren't equal so what two numbers could you put x and y equal two so that the two expressions are the same?